OpenStudy (anonymous):

Identify the type of conic section that has the equation 9x2+ 16y2 = 144 and identify its domain and range. ellipse Domain: {–3 ≤ x ≤ 3} Range: {–4 ≤ y ≤ 4} ellipse Domain: {–4 ≤ x ≤ 4} Range: {–3 ≤ y ≤ 3} hyperbola Domain: all real numbers Range: {y ≤ –4 or y ≥ 4} hyperbola Domain: {x ≤ –4 or x ≥ 4} Range: all real numbers

OpenStudy (anonymous):

1. Identify the type of conic section that has the equation 9x2+ 16y2 = 144 and identify its domain and range. (1 point) (0 pts) ellipse Domain: {–3 ≤ x ≤ 3} Range: {–4 ≤ y ≤ 4} (1 pt) ellipse Domain: {–4 ≤ x ≤ 4} Range: {–3 ≤ y ≤ 3} (0 pts) hyperbola Domain: all real numbers Range: {y ≤ –4 or y ≥ 4} (0 pts) hyperbola Domain: {x ≤ –4 or x ≥ 4} Range: all real numbers 1 /1 point 2. Use the graph to write an equation for the parabola. (1 point) (0 pts) (1 pt) (0 pts) (0 pts) y = 2x2 1 /1 point 3. Find an equation of a parabola with a vertex at the origin and directrix y = –1.5. (1 point) (0 pts) x = –y2 (0 pts) x = y2 (0 pts) y = –x2 (1 pt) y = x2 1 /1 point 4. The reflecting dish of a parabolic microphone has a cross-section in the shape of a parabola. The microphone itself is placed the focus of the parabola. If the parabola is 20 inches wide and 5 inches deep, how far from the vertex should the microphone be placed? (1 point) (0 pts) 4 inches (1 pt) 5 inches (0 pts) 8 inches (0 pts) 10 inches 1 /1 point 5. What is the equation of the parabola with vertex (2, 4) and focus (5, 4)? (1 point) (0 pts) y = (x – 4)2 +2 (0 pts) y =(x – 4)2 +2 (0 pts) x = (y – 4)2 +2 (1 pt) x =(y – 4)2 +2 1 /1 point 6. Write an equation of a circle with the given center and radius. center (2, 4) and radius 5 (1 point) (0 pts) (x – 2)2 + (y + 4)2 = 5 (0 pts) (x + 2)2 + (y – 4)2 = 5 (0 pts) (x + 2)2 + (y – 4)2 = 25 (1 pt) (x – 2)2 + (y – 4)2 = 25 1 /1 point 7. Write an equation for the translation of x2 + y2 = 25 by 8 units left and 4 units down. (1 point) (0 pts) (x + 8)2 + (y – 4)2 = 25 (1 pt) (x + 8)2 + (y + 4)2 = 25 (0 pts) (x – 8)2 + (y + 4)2 = 25 (0 pts) (x – 8)2 + (y - 4)2 = 25 0 /1 point 8. Write an equation in standard form for the circle. (1 point) (1 pt) (x + 3)2 + (y + 6)2 = 16 (0 pts) (x – 3)2 + (y + 6)2 = 16 (0 pts) (x + 3)2 + (y – 6)2 = 16 (0 pts) (x – 3)2 + (y – 6)2 = 16 1 /1 point 9. What is the center and radius of the circle with the given equation? x2 + y2 + 14x – 12y = –69 (1 point) (0 pts) center at (–7, 6); radius 16 (0 pts) center at (7, –6); radius 4 (0 pts) center at (7, –6); radius 16 (1 pt) center at (–7, 6); radius 4 0 /1 point 10. An ellipse has foci (0, ±3) and vertices (0, ±4). What is the eccentricity of the ellipse? (1 point) (0 pts) 0 (0 pts) 0.25 (1 pt) 0.75 (0 pts) 1.00 1 /1 point 11. Write an equation of the ellipse with foci at (0, ±11) and vertices at (0, ±12). (1 point) (0 pts) (0 pts) (0 pts) (1 pt) 1 /1 point 12. What are the foci of the ellipse given by the equation 100x2 + 64y2 = 6,400? (1 point) (1 pt) (0, ±6) (0 pts) (±6, 0) (0 pts) (0, ±36) (0 pts) (±36, 0) 1 /1 point 13. How far apart are the foci of an ellipse with a major axis of 26 feet and a minor axis of 24 feet? (1 point) (0 pts) 4 feet (0 pts) 5 feet (1 pt) 10 feet (0 pts) 20 feet 1 /1 point 14. Graph the conic section. 9x2 – 36y2 = 324 (1 point) (1 pt) (0 pts) (0 pts) (0 pts) 1 /1 point 15. What are the foci of the hyperbola with equation 16y2 – 9x2 = 144? Graph the hyperbola. (1 point) (0 pts) foci: (0, ±5) (0 pts) foci: (±5, 0) (0 pts) foci: (±5, 0) (1 pt) foci: (0, ±5)

OpenStudy (anonymous):

@ xojesse00 OMG thank you so much :D

OpenStudy (confluxepic):

All the questions are modified in some way @~ @confluxepic

OpenStudy (omarbirjas):

@confluxepic you're right, or maybe they're just old......

OpenStudy (anonymous):

Darnit. . .

OpenStudy (happpily):

@-etherealebullience are you in connections?

OpenStudy (anonymous):

@Happpily... I am. Did you finish this test?

OpenStudy (happpily):

@kstate i did :D

OpenStudy (anonymous):

Can you help me?

OpenStudy (anonymous):

@Happpily ^

OpenStudy (happpily):

@Kstate i can msg me :)

OpenStudy (anonymous):

i emailed you

OpenStudy (confluxepic):

Okay. I commented on this extremely old post 4 months ago. You guys should read the rules of OpenStudy. http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions OpenStudy has an anti-cheating policy. If you have a question then ask it and try to learn. Hopefully the person who will answer your question will guide you to the answer so that you learn throughout the process. Have fun on OpenStudy! :)

OpenStudy (anonymous):

@Happpily

OpenStudy (confluxepic):

@-EtherealEbullience @Happpily @Kstate

OpenStudy (happpily):

@confluxepic i know the rules thanks for the reminder ... never actually implied we were going to cheat but okay....

OpenStudy (anonymous):

@Happpily can you help me out with this?